OpenStudy (anonymous):

help? If f(x) = x - 4, what does f(3) represent? The value of (x - 4) when x = 3 The value if x when (x - 4) = 3 The value of 3f(x) when x = 3 The value of f(x) when (x - 4) = 3

OpenStudy (anonymous):

The value of (x-4) when x=3 is -1 The value of x when (x-4)=3 is x= 7 The value of 3f(x) when x=3 is 9 The value of f(x) when (x-4)=3 is 7

OpenStudy (anonymous):

for the fist one: you replace x with 3 to make (3-4) 2. (x-4)=3 add 4 to the right side to make x=7 3. if x=3 then you multiply 3 and 3 to get 9 4. you do the same thing as 2

OpenStudy (anonymous):

i know, but its asking me which answer to choose? @lmeithof

OpenStudy (anonymous):

Oops sorry :( the answer would be the first one because when you say f(3), you are keeping the equation the way it is, just turning x into 3

OpenStudy (anonymous):

@lonelyflowers

OpenStudy (anonymous):

thank you :) @lmeithof

OpenStudy (anonymous):

you're welcome :)